# Gravitational wave probes of particle dark matter: a review

### v3

@@ III.1.2 Excess power time-frequency method
Cross-correlation is able to capture the phase information in each FFT; however, at the moment the software has not yet been implemented to change $T_{\text{FFT}}$ to match $T_{\text{coh}}$, which would result in optimal sensitivity to each DM mass. Thus, another method Miller et al. (2021) was developed that can vary $T_{\text{FFT}}$ as a function of frequency, and that is more robust against noise disturbances. It relies on creating time-frequency spectrograms in which frequencies in each FFT are only kept if their powers are (1) above a given threshold and (2) local maxima. After applying these cuts, a time-frequency “peakmap” is created, which is a collection of ones that indicates particular time/frequency points at which the aforementioned two conditions are met. A peakmap can be created every Hz, and $T_{\text{FFT}}$ can be varied to match $T_{\text{coh}}$ in each 1 Hz window. Because $T_{\text{FFT}}\sim T_{\text{coh}}$, the signal is expected to be sinusoidal. Thus, we can sum, in each frequency bin, the ones that are present, in essence creating a histogram. On this histogram, at each frequency, we compute a detection statistic called the “critical ratio”:
@@
where $n$ is the number of peaks at a given frequency, and $\mu$ and $\sigma$ are the mean and standard deviation of the number of peaks in the histogram. The CR, like the SNR, follows a normal distribution. The minimum detectable strain amplitude of this method is:
$h_{\text{0,min}}$ $\approx\frac{\mathcal{G}}{N_{\text{FFT}}^{1/4}\theta^{1/2}_{\rm thr}}\sqrt{\frac{S_{n}(f)}{T_{\text{FFT,max}}}}$
×$\times\left(\frac{p_{0}(1-p_{0})}{p_{1}^{2}}\right)^{1/4}\sqrt{CR_{\text{thr}}-\sqrt{2}\mathrm{erfc}^{-1}(2\Gamma)},$
$N_{\text{FFT}}$ $=\frac{T_{\text{obs}}}{T_{\text{FFT,max}}},$
$p_{0}$ $=e^{-\theta_{\text{thr}}}-e^{-2\theta_{\text{thr}}}+\frac{1}{3}e^{-3\theta_{\text{thr}}}=0.0755\text{ for $\theta_{\text{thr}}=2.5$,}$
$p_{1}$ $=e^{-\theta_{\text{thr}}}-2e^{-2\theta_{\text{thr}}}+e^{-3\theta_{\text{thr}}}\text{ }=0.0692\text{ for $\theta_{\text{thr}}=2.5$,}$ (75)
where $\mathcal{G}=[1.3,2.8,2.08]$ for scalar, vector or tensor DM interactions, respectively. The $\mathcal{O}(1)$ differences arise from averages over the spin-1 or spin-2 polarizations. $\theta_{\text{thr}}$ is a threshold on the equalized power in the time-frequency spectrogram, and $\Gamma$ is the confidence level.
@@ III.1.3 Logarithmic power spectral density (LPSD) method
To truly maximize the sensitivity towards DM signals, one must, with infinite precision, analyze chunks of data with $T_{\text{coh}}$ as the DM mass changes. LPSD Tröbs and Heinzel (2006); Vermeulen et al. (2021); Göttel et al. (2024) is an attempt to do this, which creates spectra that are logarithmically spaced in frequency. It is extremely computationally intensive, since it requires an implementation separate from the discrete FFT. The benefit of using such an algorithm is its ability to estimate power spectral densities logarithmically, thereby capturing important features of the signals that could be missed when estimating the power spectral density across a linear frequency range. Given that the ultralight dark matter causes frequency variations of $\mathcal{O}(10^{-6})$, it is worthwhile to use LPSD to capture these tiny fluctuations in frequency.
Accounting for the strain on the test masses, the loss of sensitivity due to the number of round-trips of the laser light, as well, the induced strain can also be written in terms of transfer functions between scalar DM and the interferometers:
$h(\omega)\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\cdot\left(\frac{\hbar\,\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}\,c}\right)A_{\rm cal}^{-1}(\omega)$ (76)
where $A_{\rm cal}$ is a function given in Göttel et al. (2024) that encapsulates the transfer functions between the DM and the beam splitter and test masses, as well as the thicknesses and indices of refraction of the mirrors, the mirrors’ transmissivities, and the GW/ detector transfer function. Then, $h(\omega)$ can be estimated from the data using Eq. 11 and 12 in Göttel et al. (2024), and Eq. 16 in Cowan et al. (2011), which can be mapped to constraints on the coupling constants using Eq. 76.
@@ III.1.4 Stochastic summing method
This method leverages the fact that the signal power will be stochastically spread at frequencies above the characteristic one if $T_{\text{FFT}}>T_{\text{coh}}$, and essentially sums the power at those frequencies. Afterwards, a detection statistic, like the signal-to-noise ratio, is calculated by dividing by the noise power spectral density:
$\rho(f_{c})\equiv\sum_{i}^{N_{\text{FFT}}}\sum_{f_{0}\leq f_{n}\leq f_{0}(1+\kappa^{2}v^{2}_{0}/c^{2})}\frac{4|\tilde{d}(f_{n};t_{i})|^{2}}{T_{\text{FFT}}S(f_{n};t_{i})},$ (77)
@@
where $\tilde{d}(f_{n};t_{i})$ is the FFT of $i^{\rm th}$ data chunk, $S(f_{n};t_{i})$ is the one-sided noise power spectral density at $t=t_{i}$, and $\kappa$ = 3.17 to ensure that no more than 1% of signal power is lost by restricting the inner sum. The outer sum is over all chunks of length $T_{\text{FFT}}$; the inner sum is over the frequency spread of the signal $\delta f_{v}$.
@@
The sensitivity of this method has been estimated in Nakatsuka et al. (2023) for both axions and dark photons, analytically and numerically in the case of $T_{\text{FFT}}=T_{\text{obs}}$. In the absence of a signal, assuming Gaussian noise, a detection threshold can be established as a function of the false alarm probability $\alpha$ (Eq. 47 in Nakatsuka et al. (2023)), and this threshold can be fed back into an integral over the likelihood function (Eq. 46 in Nakatsuka et al. (2023)), to compute, at a given confidence level, the minimum detectable signal amplitude $\bar{\lambda}_{\mathrm{up}}$:
$\bar{\lambda}_{\mathrm{up}}(T_{\text{obs}})=\begin{cases}(\Delta_{X,\mathrm{tot}})^{-1/2}\sqrt{\frac{\ln(\alpha)}{\ln(\beta)}-1},&\mathrm{for}~{}T_{\text{obs}}<\tau/\kappa,\\ (\Delta_{X,\mathrm{tot}})^{-1/2}\sqrt{M_{\alpha}+M_{1-\beta}}\left(\kappa T_{\text{obs}}/T_{\text{coh}}\right)^{1/4}&\mathrm{for}~{}T_{\text{obs}}\gg T_{\text{coh}}.\end{cases}$ (78)
$\Delta_{X,\text{tot}}$ encodes the deterministic part of the spectral shape of the signal, and depends on the relative motion of the DM velocity with respect to the interferometer arms. $\beta$ is the detection probability. In principle, Eq. 11 and 17 of Nakatsuka et al. (2023) give these relative motion terms; in practice, however, conservative ($\vec{v}=[0,0,v_{0}]$) and optimistic ($\vec{v}=[v_{0},0,0]]$) choices of the direction of the DM velocity with respect to the interferometers result in values of [0.38,0.81], respectively (see Eq. 52 in Nakatsuka et al. (2023)). $M_{\chi}$ represents the relation between the peak width and the area of the Gaussian distribution defined by

### v4

@@ III.1.2 Excess power method
Cross-correlation is able to capture the phase information in each FFT; however, at the moment the software has not yet been implemented to change $T_{\text{FFT}}$ to match $T_{\text{coh}}$, which would result in optimal sensitivity to each DM mass. Thus, another method Miller et al. (2021a) was developed that can vary $T_{\text{FFT}}$ as a function of frequency, and that is more robust against noise disturbances. It relies on creating time-frequency spectrograms in which frequencies in each FFT are only kept if their equalized powers are (1) above a given threshold $\theta_{\rm thr}=2.5$ and (2) local maxima. The equalized power is computed at each time and frequency bin by calculating the ratio $R_{ij}$ of the square modulus of the FFT with a running median estimation of the power spectral density, as in Section III.1.1:
$R_{ij}=\frac{|FFT|_{I,ij}^{2}}{P_{I,ij}}$ (69)
which, on average, takes on $\mathcal{O}(1)$ values. After applying these cuts, a time-frequency “peakmap” is created, which is a collection of ones that indicates particular time/frequency points at which the aforementioned two conditions are met (the value $R_{ij}$ does not enter into any subsequent part of the analysis). A peakmap can be created every Hz, and $T_{\text{FFT}}$ can be varied to match $T_{\text{coh}}$ in each 1 Hz window. Because $T_{\text{FFT}}\sim T_{\text{coh}}$, the signal is expected to be sinusoidal. Thus, we can sum, in each frequency bin, the ones that are present, in essence creating a histogram. On this histogram, at each frequency, we compute a detection statistic called the “critical ratio” $CR$:
@@
where $n$ is the number of peaks at a given frequency, and $\mu$ and $\sigma$ are the mean and standard deviation of the number of peaks in the histogram. The $CR$, like the SNR, follows a normal distribution. Using the $CR$ and assuming Gaussian noise, the minimum strain amplitude of a sinusoidal signal can be derived that would, in a frequentest interpretation, produce a detectable signal in a fraction $\geq\Gamma$ of a repeated number of experiments Astone et al. (2014); Miller et al. (2021a):
$h_{\text{0,min}}$ $\approx\frac{\mathcal{G}}{T_{\text{obs}}^{1/4}T_{\text{FFT}}^{1/4}}\sqrt{\frac{P_{I}(f)}{\theta_{\text{thr}}}}$
$\times\left(\frac{p_{0}(1-p_{0})}{p_{1}^{2}}\right)^{1/4}\sqrt{CR_{\text{thr}}-\sqrt{2}\mathrm{erfc}^{-1}(2\Gamma)},$
$p_{0}$ $=e^{-\theta_{\text{thr}}}-e^{-2\theta_{\text{thr}}}+\frac{1}{3}e^{-3\theta_{\text{thr}}},$
$p_{1}$ $=e^{-\theta_{\text{thr}}}-2e^{-2\theta_{\text{thr}}}+e^{-3\theta_{\text{thr}}}$ (71)
where $\mathcal{G}$ depends on whether DM is scalar, vector or tensor:
$\mathcal{G}$ $=\sqrt{\frac{2\pi}{2.4308}}\frac{\sqrt{2}}{3}\times\begin{cases}1,\text{\quad\, dilatons}\\ \sqrt{\frac{9}{2}},\text{ dark photons}\\ \sqrt{\frac{5}{2}},\text{ tensor bosons}\\ \frac{3}{\sqrt{2}}\frac{5}{2},\text{ CWs}\end{cases}$ (72)
$=\begin{cases}1.31,\text{ \, dilatons}\\ 2.78,\text{ dark photons}\\ 2.08,\text{ tensor bosons}\\ 4.02,\text{ CWs}\end{cases}$ (73)
The $\mathcal{O}(1)$ differences arise from averages over the spin-1 or spin-2 polarizations, and are valid for ground-based L-shaped GW interferometers (for triangular shapes, the averages will be slightly different). The $\frac{\sqrt{2}}{3}$ factor comes from an average over the DM Maxwell-Boltzmann distributed velocities as discussed in Section II.4, while the $\frac{2\pi}{2.4308}$ factor comes from convolving a sinusoid with a rectangular window function when computing the sensitivity of GW interferometers to a sinusoidal signal Astone et al. (2014).
@@ III.1.3 Logarithmic power spectral density (LPSD) method
To maximize sensitivity to ultralight dark-matter signals, it is important to let $T_{\text{FFT}}$ match $T_{\text{coh}}$ as the DM mass varies. The logarithmic power spectral density (LPSD) method Tröbs and Heinzel (2006); Vermeulen et al. (2021); Göttel et al. (2024) implements this idea by constructing spectra that are logarithmically spaced in frequency, in contrast to the excess power method in Section III.1.2 that use linearly spaced spectra every 1 Hz. While this approach is computationally intensive due to the need for a separate implementation from the standard discrete FFT, it captures subtle features in the data that could otherwise be missed. In particular, since ultralight dark matter induces fractional frequency variations on the order of $\mathcal{O}(10^{-6})$, a logarithmic spectral estimate is well-suited to resolving such tiny modulations.
Once the LPSD is computed, it can be combined with the interferometer response to convert measured strains into constraints on DM couplings. Accounting for the strain induced on the test masses and beam splitter (see Section II.3.1), the expected signal can be written as
$h(\omega)\approx\left(\frac{1}{\Lambda_{\gamma}}+\frac{1}{\Lambda_{e}}\right)\cdot\left(\frac{\sqrt{2\,\rho_{\mathrm{DM}}}}{m_{\rm DM}}\right)A_{\rm cal}^{-1}(\omega)\,$, (74)
where $A_{\rm cal}(\omega)$ encodes the full interferometer transfer function. This function accounts for the response of the beam splitter and test masses to the DM signal, which is amplified by the difference in mirror thicknesses as discussed in Section II.3.1, mirror transmissivities, and the overall canonical GW detector transfer function.
By applying Eq. 74 to the logarithmically estimated spectra, one can extract an estimate of the strain $h(\omega)$ from the data. This estimate can then be mapped to constraints on the coupling constants $\Lambda_{\gamma}$ and $\Lambda_{e}$, providing a direct connection between the interferometer measurement and the underlying ultralight dark matter model Göttel et al. (2024); Cowan et al. (2011). In this way, the LPSD method offers a way to detect logarithmically-spaced frequency modulations that might be missed by linear spectral analyses.
@@ III.1.4 Stochastic summing method
The stochastic summing method Nakatsuka et al. (2023) exploits the fact that when $T_{\text{FFT}}>T_{\text{coh}}$, the signal power from ultralight dark matter is spread stochastically over a range of frequencies rather than concentrated at a single frequency. Instead of searching for a narrow peak, this method sums the power across all frequency bins where the signal is expected, producing a detection statistic that captures the total available signal power.
$\rho(f_{c})\equiv\sum_{i}^{N_{\text{FFT}}}\sum_{f_{0}\leq f_{n}\leq f_{0}(1+\kappa^{2}v^{2}_{0})}\frac{4|\tilde{d}(f_{n};t_{i})|^{2}}{T_{\text{FFT}}S(f_{n};t_{i})},$ (75)
@@
where $\tilde{d}(f_{n};t_{i})$ is the FFT of $i^{\rm th}$ data chunk, $S(f_{n};t_{i})$ is the one-sided noise power spectral density at $t=t_{i}$, and $\kappa$ = 3.17 is chosen to ensure that no more than 1% of signal power is lost by limiting the frequencies summed over. The outer sum is over all chunks of length $T_{\text{FFT}}$; the inner sum is over the frequency spread of the signal $\delta f_{v}$.
@@
The sensitivity of the stochastic summing method has been studied for both axions and dark photons Nakatsuka et al. (2023), using a combination of analytical estimates and numerical simulations. In the absence of a signal and assuming Gaussian noise, a detection threshold can be established based on a chosen false-alarm probability. This threshold can then be used to determine, at a given confidence level, the minimum detectable signal amplitude, similar to Eq. 71 in Section III.1.2.
The exact sensitivity depends on the spectral shape of the signal, which in turn is influenced by the relative motion of the DM with respect to the interferometer. Conservative and optimistic assumptions about the direction of the DM velocity lead to $\mathcal{O}(1)$ variations in the sensitivity, which are accounted for in this method Nakatsuka et al. (2023).
Once the minimum detectable strain amplitude is determined, it can be translated into upper limits on the relevant coupling constants for each type of dark matter interaction, similarly to what has been discussed in Sections III.1.1 and III.1.2. These include axion-photon couplings as well as scalar or vector couplings that induce displacements or charge effects on the interferometer mirrors. In practice, because data are typically divided into segments with $T_{\text{FFT}}\sim T_{\text{coh}}$, the analytic likelihood approach can become unstable. Consequently, numerical simulations are used to estimate the sensitivity and derive robust upper limits, as done in a search on O3 KAGRA data Abac et al. (2024), which will be discussed in Section III.2.
III.1.5 Distinguishing amongst dark-matter models
Only a few works Miller et al. (2022b); Manita et al. (2024) have explored ways to determine whether a particular type of DM interacted with GW detectors. In Miller et al. (2022b), the authors employ the Wiener filter Wiener et al. (1964) to compute the cross power across different detector pairs during follow-up stages, using different $T_{\text{FFT}}$ to exploit the fact that different types of DM couple differently to different detector baselines.
In Wiener filtering, the goal is to find a waveform that, when subtracted from the data, leaves only noise. The residual quantifies the mismatch between the data, $h(t)+n(t)$, and a candidate signal, $h(t)$. Low residuals indicate that the candidate waveform closely matches the data, whereas high residuals indicate poor agreement.
Miller et al. (2022b) showed that the Wiener filter produces small residuals when the data contain a signal consistent with the model and large residuals when the wrong signal is applied. This property allowed the method to confidently veto spurious candidates in real O3 data. Conversely, this means that the method has the potential to confirm the detection of a certain type of DM interaction signal. This is in contrast to other methods and previous searches Guo et al. (2019); Miller et al. (2021a); Abbott et al. (2022c), which focused mostly on rejecting outliers that were shown to be compatabile with noise.
After the work of Miller et al. (2022b), Ref. Manita et al. (2024) uses standard cross correlation to argue that the overlap reduction function can be used as way to distinguish amongst different spins of DM particles, since the cross power will be different. In particular, Ref. Manita et al. (2024) how the cross-correlation SNR would change as a function of the type of DM and which baseline was considered. In both cases, cross power is important, since the individual power spectra from one detector are indistinguishable for all types of ultralight dark matter, which arises because the phase evolutions for the considered types of ultralight dark matter here are the same.
III.2 Ground- and space-based interferometers
Searches for DM particles interacting with the LIGO, Virgo, KAGRA and GEO600 interferometers have been performed using data from the most recent observing run, O3, resulting in constraints on dilatons and dark photons Abbott et al. (2022c); Vermeulen et al. (2021); Göttel et al. (2024); Abac et al. (2024). Furthermore, LISA Pathfinder data were also used to set constraints on dark photon DM Miller and Mendes (2023); Frerick et al. (2024). The theoretical framework for each particle is described in Section II.3 and Section II.4, respectively. In this subsection, we summarize key constraints from existing ground-based interferometers on these two particles.
III.2.1 LIGO, Virgo and GEO600
In Ref. Abbott et al. (2022c), two complementary methods were employed to search for spin-1 ultralight dark matter that can couple to baryons in the test masses of LIGO in O3 data – cross-correlation and excess power, both of which are described in Section III.1.1 and Section III.1.2, respectively. Cross-correlation benefits from utilizing the phase information of the signal, something that is lost when using the time-frequency excess power method. On the other hand, the excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every one-Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
Both searches returned no significant DM candidate signals; thus, 95% confidence-levle upper limits were set on the coupling strength $\epsilon^{2}$ of dark photons to baryons for both methods, as shown in Fig. 4.. We can see, across a broad frequency range, that results from GW interferometers surpass by orders of magnitude constraints from Eöt-Wash and the MICROSCOPE experiments. These upper limits were derived for cross-correlation and excess power following the procedure outlined in Section III.1.1 and Section III.1.2, respectively. Both methods employed the Feldman-Cousins approach Feldman and Cousins (1998) to set upper limits, which is robust against non-Gaussian noise and has been shown to produce conservative upper limits with respect to those that would have been obtained through simulations Miller et al. (2021a).
Figure 4: Adapted from Abbott et al. (2022c). Upper limits obtained from analyzing LIGO O3 data on the square of the coupling of dark photons to baryons $U(1)_{\rm B}$ in the LIGO mirrors. The physics behind this form of DM has been discussed in Section II.4. The limits derived from each of the methods discussed in Sections III.1.1 and III.1.2 are show in red (cross-correlation) and cyan (excess power), respectively. MICROSCOPE Bergé et al. (2018) and Eöt-Wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger et al. (2008). To produce limits on the square of the dark photon/baryon-lepton coupling, $U(1)_{\rm B-L}$ in the LIGO mirrors, these limits should be multiplied by four.
While cross-correlation and excess power analyses were used for dark photon searches, the LPSD technique (described in Section III.1.3) was employed in Ref. Göttel et al. (2024) to search for scalar dilaton dark matter in the LIGO O3 dataset. This analysis placed strong limits on the coupling of scalar dark matter to electrons and photons. In Fig. 5, we can see, in blue, the constraints that have been derived, which supersede, at least at low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen et al. (2021). The enhancement comes from the improved low-frequency sensitivity of LIGO relative to GEO600, as well as incorporating the interaction of scalar DM with the mirrors, as well as the beam splitter, into the search. However, at high frequencies, we can see that GEO600 still outperforms LIGO, primarily because of sophisticated quantum technologies that were employed in its design. These upper limits were derived by following the procedure outlined in Section III.1.3 and Section III.1.2.
Figure 5: Taken from Göttel et al. (2024). 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, the coupling of dilaton DM to electrons or photons, as a function of mass and frequency from LIGO O3 data, which would have cause time-dependent oscillations of the sizes and indices of refraction of the beamsplitter and the LIGO mirrors. The physics behind this form of DM has been discussed in Section II.3.1.a) and b) show these constraints compared to existing ones on $\Lambda_{e}$ and $\Lambda_{\gamma}$, respectively. The results from the LIGO O3 search are shown by the thick blue line, constraints from direct experimental searches for DM Vermeulen et al. (2021); Aiello et al. (2022); Aharony et al. (2021); Savalle et al. (2021); Kennedy et al. (2020); Antypas et al. (2019, 2021); Tretiak et al. (2022); Oswald et al. (2022); Campbell et al. (2021); Beloy et al. (2021); Zhang et al. (2023a); Fukusumi et al. (2023); Sherrill et al. (2023) are shown in thin grey, and constraints from searches for “fifth forces” Bergé et al. (2022); Hees et al. (2018) are depicted by the dashed red lines.
III.2.2 KAGRA
KAGRA offers unique opportunities to search for DM. In particular, the different materials used in KAGRA’s mirrors result in significantly different charge-to-mass ratios for baryon-lepton number, enhancing the sensitivity to vector DM coupling to this quantum number Michimura et al. (2020). While LIGO and Virgo currently set the strongest limits on dark photon coupling to baryon-lepton number – obtainable from the results in Fig. 4 by multiplying the limits on baryon number coupling by four Abbott et al. (2022c) – this difference in mirror materials means that once KAGRA reaches comparable sensitivity levels, it will outperform LIGO in constraining such couplings Michimura et al. (2020).
This enhancement is particularly apparent in KAGRA’s auxiliary length control channels, PRCL and MICH, which record displacements of mirrors made from both fused silica (beam splitter and power recycling mirrors) and cryogenic sapphire (test masses). The baryon-lepton charge-to-mass ratios differ by approximately $9\times 10^{-3}$ between these materials, compared to only about $4\times 10^{-5}$ for baryon number alone, significantly amplifying the DM interaction signal in these channels – see Eq. 54.
The raw outputs of PRCL and MICH correspond to time-varying displacements of the form Abac et al. (2024):
$\delta L_{\rm MICH}$ $=\delta(l_{\rm x}-l_{\rm y}),$ (76)
$\delta L_{\rm PRCL}$ $=\delta\left[\frac{l_{\rm x}+l_{\rm y}}{2}+l_{\rm p}\right],$ (77)
where $l_{x}=26.7\,\text{m}$ and $l_{y}=23.3\,\text{m}$ are the distances between the beam splitter and the input mirrors, and $l_{p}=41.6\,\text{m}$ is the length of the power recycling cavity. All these lengths are much shorter than the DM coherence length $L_{\text{coh}}$. See Fig. 1 of Abac et al. (2024) for a schematic of the interferometer.
No significant signals were found in this analysis of KAGRA O3 data Abac et al. (2024), and upper limits on the coupling strength were derived following the procedure outlined in Section III.1.4. These limits, shown in Fig. 6 for each channel, exhibit many narrow-band noise artifacts that limit sensitivity at low frequencies. Although not yet competitive with LIGO constraints, future KAGRA searches are expected to provide powerful limits on this type of DM due to the pronounced difference in baryon-lepton charge-to-mass ratios among its mirrors Michimura et al. (2020).
Figure 6: Taken from Abac et al. (2024). 95% confidence-level upper limit on the strength of the coupling between dark photons and baryon-lepton number using KAGRA O3 data and the method described in Section III.1.4, derived from two particular interferometer channels: MICH (blue line) and PRCL (orange line). The physics behind this form of DM has been discussed in Section II.4. Because the KAGRA input mirrors, end mirrors, and power recycling mirrors are made of different material, the coupling strength of dark photons to baryon-lepton number is enhanced (Eq. 54), and appears more strongly in these particular channels than in the ordinary strain one. In the low-mass (low-frequency) range, many narrow-band noise disturbances of unknown origin are seen.
III.2.3 LISA Pathfinder